OpenStudy (anonymous):

t= -11+- (sq root) 81 /4 my teacher wants simplified form so would it just be t= -11+-9/4?

OpenStudy (nirmalnema):

-11 + - 9/2 (-22(-+) 9 ) / 2

OpenStudy (nirmalnema):

\[\sqrt{81/4} or \sqrt{81}/4\]

OpenStudy (anonymous):

\[\sqrt{81}/4=4.5\]

OpenStudy (anonymous):

Actually the final answers were t=0, -1/2, -5